Please help me solve this problem in steps y= -7x - 4 (where y= -2)
Thank You

Answer by checkley77(12844)

y= -7x - 4 (where y= -2)
-2=-7x-4
-7x=-2+4
-7x=2
x=2/-7
x=-2/7 answer.